OpenStudy (anonymous):

a population of bacteria is growing acording to the equation p(t)=1200e^0.17t, with t measured in years. estimate when the population will exceed 3443

OpenStudy (anonymous):

\[3343\leq 1200e^{.07t}\] \[\frac{3343}{1200}\leq e^{.17t}\] \[\ln(\frac{3343}{1200})\leq .17t\] \[\frac{\ln(\frac{3343}{1200})}{.17}\leq t\]

OpenStudy (anonymous):

what would be the final answer